# Probing the radiation-dominated regime of laser-plasma interaction in multi-beam configurations of petawatt lasers

### v1

@@ II.2 Numerical methods
In the simulations, 64 macroions and macroelectrons per cell were used at the spatial resolution $\Delta x,\Delta y,\Delta z=\lambda/40.$ Boundary conditions for the fields on the left plane of the computational domain, $x=-5\lambda$, are given by (3). Otherwise, the conditions for the fields and particles are: absorption in the propagation direction and periodicity in the lateral directions. The size of the simulation box $[35\times 25\times 25]\lambda^{3}$ is chosen sufficiently large to rule out any boundary effects on the particles’ and fields’ distributions in the central part of the simulation domain.
@@
Here, $r_{e}=e^{2}/mc^{2}\approx 2.8\times 10^{-13}$cm is the classical electron radius, $\mathbf{v}$ is the electron velocity, $\mathbf{F}_{L}=\left(\mathbf{E}+\mathbf{v}/c\times\mathbf{B}\right)$ nd $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic gamma-factor. This is a reduced expression for the RRF, compared to its full Landau-Lifshitz form [4]. It follows from the latter after omitting terms (small in the considered case) containing time and space derivatives of the electromagnetic fields. Note that here we treat radiation reaction classically, which is amply sufficient for the considered parameters of the laser pulses. For account of quantum effects in descriptions of the RRF and in particular of the IFE, see reviews [10, 2] and papers [32, 33, 16, 34]. The codes use a numerical algorithm based on the assumption that the acceleration of particles in their rest frame is dominated by the Lorentz force (see Ref. [35] for detail). The numerical implementation allows adding radiation reaction effects to any PIC code based on the standard Boris pusher algorithm for the calculation of the particles acceleration with small computational costs.
Inclusion of dissipative radiation losses through the RRF in a PIC code is justified if the following conditions are met: 1) characteristic frequencies of synchrotron-like radiation emitted by the plasma electrons are much higher than the highest frequency that can be resolved on the numerical grid; 2) radiation at such frequencies is incoherent; 3) the plasma is transparent for electromagnetic fields of these frequencies. For the present calculations, these requirements are well satisfied (see estimates in Refs.[14, 16]). At lower though still ultra-relativistic intensities, radiation responsible for the RRF may become partially coherent, which requires to revisit the description of radiation reaction effects. For this specific case, see recent publications [36, 37].
@@ III Results and discussion
Figure 2: Spatial distributions in $B_{x}/B_{0}$ as functions of $x$ and $r=\sqrt{y^{2}+z^{2}}$ at $t=35T_{L}$ for the cases of two (a,c) and four (b,d,f) beams with the angle of convergence (see Fig. 1 and text around for explanations) $2\theta=5^{\rm o}$ (a), (b); $2\theta=10^{\rm o}$ (c), (d) and $2\theta=20^{\rm o}$ (f). The single beam power was fixed by 16PW corresponding to $a_{0}=200$. For reference, panel (e) shows the distribution for a single CP beam with power $P=64$PW equal to the total power of the four beams. Panel (c) gives the distribution for the same case as (a), but assuming 32PW $(a_{0}=280)$ in each beam instead of 16 PW. Note different color scales in panels (a,c) and (b,d,e,f) correspondingly. The lines of constant magnetic field $B_{x}=[-15,-10,-5]B_{0}$ are displayed white.
@@
    When the IFE is present for collinear beams, it survives for the angles of convergence up to $2\theta\simeq 10^{\rm o}$. For $2\theta\simeq 20^{\rm o}$ the magnetic field drops down considerably, both in the amplitude and the size of space where the magnetic field is present. These features are similar for 2 and 4 beams. For the 2-beam scheme a similar behavior was reported in a brief publication [27].
@@
    The difference between the 2- and 4-beam schemes is evident for the fixed power of a single beam. The distributions of Fig. 2(a,c) show that for the 2-beam scheme the threshold power, when some signatures of the magnetic field generated through the IFE can be expected, is $\simeq 30$PW, well above the presently achieved level of 5PW and the soon expected that of 15PW. Instead, the 4-beam scheme should work reliably at the 15PW power. For the scale $B_{x}\in[+10,-10]B_{0}$ of panels (a) and (c) no magnetic field is visible when two 15PW beams are applied. Reduction of the interval reveals some fields $B_{x}$ which however lie beyond the reliable signal level of the present PIC simulations.
@@
    The quasi-static longitudinal magnetic field $B_{x}$ vanishes in correlation with the decrease in the longitudinal projection of the field angular momentum (see Fig. 1), which agrees well with macroscopic picture of the IFE [14] based on the irreversible transfer of angular momentum from the field to the particles. However, one can notice from the comparison of Fig. 1(c) and Fig. 2 (d,f) that the magnetic field of negative sign survives even when the sign of $M_{x}$ is changed. An explanation of this extended survival of the IFE with the angle between the beams increasing stems from the distributions shown in Fig. 1(f,g) and (h-k). It is clear from these plots that the density of $m_{x}$ remains positive in the central part of the interaction volume, up to $r\approx 3\lambda$, while the electron current generating the field is concentrated near $r\approx 2\lambda$. The lateral electric field although oscillates at $y=4\lambda$, still contains a significant fraction of a CP component. The periphery of the interaction volume, which contributes a lot into the net angular momentum $M_{x}$ still plays a minor role in the excitation of the electron current.
@@
Secondly, for multi-PW laser sources with sub-kJ pulse energies and low repetition rates, it is hard to expect the beams to be well synchronized in phase and amplitude. For such laser systems under development, the option of fully incoherent beams is also considered as one of realistic scenarios. Thus it is of high relevance for the future experiments to make a step out of the idealistic assumption of fully coherent in-phase beams and to study sensibility of the effect to the fluctuations of relative phases and amplitudes of the beams. Fig. 3 presents the magnetic field distribution $B_{x}(r,x)/B_{0}$ recorded at $t=35T_{L}$ for the 4-beam case with $2\theta=5^{\rm o}$ and $2\theta=10^{\rm o}$. These cases were chosen basing on the analysis of the distributions present in Fig. 2. Panels (a-f) correspond to regular phase shifts between the pairs of beams with orthogonal polarizations, i.e. between the two blue and the two red beams in Fig.1(b), while (g) and (h) show the distributions generated by beams with randomly shifted phases and amplitudes, correspondingly.
@@
Figure 3: Magnetic field distribution for the 4-pulse scheme with $2\theta=5^{\rm o}$ (a,c,e) and $2\theta=10^{\rm o}$ (b,d,f) calculated at $t=35T_{L}$ for the phase shift $\Delta\varphi=0.0$ (a,b), $\Delta\varphi=0.2\pi$ (c,d) and $\Delta\varphi=0.4\pi$ (e,f). The field amplitude $a_{0}=250$ in each beam corresponds to the peak intensity $J\approx 1.4\cdot 10^{23}{\rm W}/{\rm cm}^{2}$ and power $P\approx 20$PW per beam. Panel (g) shows the distribution resulting from random relative phase shifts $0.075\pi,-0.018\pi,0.05\pi$; panel (h) – the distribution for randomly chosen amplitudes $a_{0}=184,242,162$ and $239$. The lines of constant magnetic field $B_{x}=[-15,-10,-5]B_{0}$ are displayed white.
@@
As expected, the efficiency of the magnetic field generation drops down with the phase shift increasing. This suppression can be qualitatively demonstrated by plotting contours of the constant magnetic field for different phase shifts. The result is shown in Fig. 4. It is important to note that the effect survives in a broad interval of relative phases. Comparison of panels (c), (d) (e) and (f) in Fig. 3 shows that the magnetic field remains visible at relative phases filling more than 50% of the interval $(0,\pi/2)$. The distributions in Figs.3 and 4 were calculated for higher field amplitudes $a_{0}=250$ and $a_{0}=500$. In the latter case, the 2-beam scheme was considered. A comparison of results presented in these Figures shows, in particular, that both the schemes are, on the qualitative level, equally sensitive to fluctuations of relative phases of the beams. In experiment, both the relative phases and amplitudes of the beams will fluctuate. Panels (g) and (h) of Fig.3 show the distributions for randomly chosen phase shifts and amplitudes. These results also suggest that the scheme is acceptably stable with respect to these parametric uncertainties.
@@ IV Conclusion
In conclusion, we have studied the possibility to reach the radiation-dominated regime of the laser plasma interaction and, in particular, to observe the Inverse Faraday effect driven by radiation reaction in a multi-beam scheme. Our numerical results show that application of four beams partially synchronized in phase and amplitude lowers the single beam power required for reliable detection of the quasi-static longitudinal magnetic field excited through the IFE to 15 PW, which looks a realistic value for several forthcoming laser facilities. Most importantly, the effect is shown rather robust with respect to fluctuations of relative phases and amplitudes. This enhances the possibility of observation at low repetition rates typical for multi-PW laser sources. At the future XCELS facility [22] installations with up to 12 beams are considered. From one side, increasing the number of beams will obviously further lower the power threshold required for IFE observation. From the other side, taking into account the constraints on the angle $\theta$ and the relative phase shifts we formulated in this paper, application of $\sim 10$ or more beams may appear highly problematic from the technical viewpoint. We conclude that the 4-beam scheme at the 10–15PW level is optimal for the search of the giant magnetic field excitation in the radiation-dominant regime.

### v2

@@ II.2 Numerical methods
In the simulations, 64 macroions and macroelectrons per cell were used at the spatial resolution $\Delta x=\Delta y=\Delta z=~{}\lambda/40.$ Boundary conditions for the fields on the left plane of the computational domain, $x=-5\lambda$, are given by (3). Otherwise, the conditions for the fields and particles are: absorption in the propagation direction and periodicity in the lateral directions. The size of the simulation box $[35\times 25\times 25]\lambda^{3}$ is chosen sufficiently large to rule out any boundary effects on the particles’ and fields’ distributions in the central part of the simulation domain.
@@
Here, $r_{e}=e^{2}/mc^{2}\approx 2.8\cdot 10^{-13}$cm is the classical electron radius, $\mathbf{v}$ is the electron velocity, $\mathbf{F}_{L}=\left(\mathbf{E}+\mathbf{v}/c\times\mathbf{B}\right)$ and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$ is the relativistic gamma-factor. This is a reduced expression for the RRF, compared to its full Landau-Lifshitz form [4]. It follows from the latter after omitting terms (small in the considered case) containing time and space derivatives of the electromagnetic fields.
Eq.(5) treats radiation reaction classically. At the considered values of the dimensionless field amplitude $a_{0}=200-400$, classical description may appear insufficient, depending on the details of interaction including its geometry. Analysis of quantum effects in photon emission in intense electromagnetic fields and and their contribution to the RRF can be found in reviews [16, 2] and papers [33, 34, 39]. In our case, light pressure and the ponderomotove forces induced by the strong laser pulse result in a considerable longitudinal plasma acceleration, leading to predominantly co-propagative interaction geometry. This in turn suppresses the effective value of the quantum parameter $\chi=E^{\prime}/E_{\rm cr}$, which is the ratio of the field amplitude in the electron’s rest frame to the critical field of quantum electrodynamics $E_{\rm cr}=m^{2}c^{3}/e\hbar=1.32\cdot 10^{16}$/cm. As a result, the interaction proceeds in the semi-classical regime, $10.1<\chi<$ 1, where quantum recoil may play a significant role, but typically does not qualitatively alter results. The impact of quantum recoil on the IFE driven by the RRF was considered in our earlier publication [22]. Following the approach explored there, we apply a semi-classical modification of the RRF force via the $g$-factor introduced by Ritus [33, 40], with the following interpolation [41]
$g(\chi)=\bigg{(}1+12\chi+31\chi^{2}+3.7\chi^{3}\bigg{)}^{-4/9}.$ (6)
The quantum parameter $\chi$ is calculated at each time step for each macroelectron by taking the values of the electromagnetic field at the macroelectron position. The quantum-corrected RRF reads $\tilde{\mathbf{F}}_{RR}=g\mathbf{F}_{RR},$ with $\mathbf{F}_{RR}$ given by Eq.(5).
The codes use a numerical algorithm based on the assumption that the acceleration of particles in their rest frame is dominated by the Lorentz force (see Ref. [42] for detail). The numerical implementation allows adding radiation reaction effects to any PIC code based on the standard Boris pusher algorithm for the calculation of the particles acceleration with small computational costs. Inclusion of dissipative radiation losses through the RRF in a PIC code is justified if the following conditions are met: 1) characteristic frequencies of synchrotron-like radiation emitted by the plasma electrons are much higher than the highest frequency that can be resolved on the numerical grid; 2) radiation at such frequencies is incoherent; 3) the plasma is transparent to electromagnetic fields of these frequencies. For the present calculations, these requirements are well satisfied (see estimates in Refs.[20, 22]). At lower though still ultra-relativistic intensities, radiation responsible for the RRF may become partially coherent, which requires to revisit the description of radiation reaction effects. For this specific case, see recent publications [43, 44].
@@ III Results and discussion
Figure 2: Spatial distributions in $B_{x}/B_{0}$ as functions of $x$ and $r=\sqrt{y^{2}+z^{2}}$ at $t=35T_{L}$ for the cases of two (a,c) and four (b,d,f-i) beams with the angle of convergence (see Fig. 1 and text around for explanations) $2\theta=5^{\rm o}$ (a), (b), (g); $2\theta=10^{\rm o}$ (c), (d), (h) and $2\theta=20^{\rm o}$ (f), (i). The single beam power was fixed by 16PW corresponding to $a_{0}=200$. For reference, panel (e) shows the distribution for a single CP beam with power $P=64$PW equal to the total power of the four beams. Panel (c) gives the distribution for the same case as (a), but assuming 32PW $(a_{0}=280)$ in each beam instead of 16 PW. Note different color scales in panels (a,c), (b,d,e,f) and (g-i) correspondingly. The lines of constant magnetic field $B_{x}=[-15,-10,-5]B_{0}$ are shown white. The inserts in frames (b,g-i) display the one-dimensional radial profiles of $B_{x}$ for the distributions shown in frames (b) at $x=5.3\lambda$ (thin line) and in frame (e) at $x=5\lambda$ (thick line). Panels (g,h,i) show the counterparts of distributions (e,b,d) calculated with the quantum $g$-factor included.
@@
    When the IFE is present for collinear beams, it survives for the angles of convergence up to $2\theta\simeq 10^{\rm o}$. For $2\theta\simeq 20^{\rm o}$ the magnetic field drops down considerably, both in the amplitude and the occupied space. These features are similar for two and four beams setup. For the two beam scheme a similar behavior was reported in a brief publication [35].
@@
    The difference between the 2- and 4-beam schemes is evident for the fixed power of a single beam. The distributions shown in Fig. 2(a,c) demonstrate that for the 2-beam scheme the threshold power, when some signatures of the magnetic field generated through the IFE can be expected, is $\simeq 30$PW, well above the presently achieved level of 5PW and the soon expected that of 15PW. Instead, the 4-beam scheme should work reliably at the 15PW power. For the scale $B_{x}\in[+10,-10]B_{0}$ of panels (a) and (c) no magnetic field is visible when two 15PW beams are applied. Reduction of the interval reveals some fields $B_{x}$ which, however, lie beyond the reliable signal level of the present PIC simulations.
@@
    The quasi-static longitudinal magnetic field $B_{x}$ vanishes in correlation with the decrease in the longitudinal projection of the field angular momentum (see Fig. 1), which agrees well with macroscopic picture of the IFE [20] based on the irreversible transfer of angular momentum from the field to the particles. However, one can notice from the comparison of Fig. 1(c) and Fig. 2 (d,f) that the magnetic field of negative sign survives even when the sign of $M_{x}$ is changed. An explanation of this extended survival of the IFE with the angle between the beams increasing stems from the distributions shown in Fig. 1(f,g) and (h-k). It is clear from these plots that the density of $m_{x}$ remains positive in the central part of the interaction volume, up to $r\approx 3\lambda$, while the electron current generating the field is concentrated near $r\approx 2\lambda$. The lateral electric field, although oscillating at $y=4\lambda$, still contains a significant fraction of a CP component. The periphery of the interaction volume, which contributes a lot into the net angular momentum $M_{x},$ plays a minor role in the excitation of the electron current.
  4. 4.
    At small values of $\theta$, the magnetic field can be even more efficiently excited than in the fully collinear geometry. This counter-intuitive behaviour of the IFE can be noticed by comparing the magnetic field distributions shown in panels (b) and (e). In the latter, we show the distribution for a single CP beam with $a_{0}=400.$ In this case, the maximal value $B_{x{\rm max}}=15.6B_{0}$ (thick line in the insert of Fig. 2b), while for the distribution generated by four LP beams with $a_{0}=200$ each and $2\theta=5^{\rm o}$ $B_{x{\rm max}}=17.3B_{0}$ (thin line in the inset). A more significant difference between the two cases is that the magnetic field occupies a considerably larger volume in the four beams setup. In Fig. 2(b) the lateral size of the space where the magnetic field exceeds $15B_{0}$ is more than three times bigger than that in Fig. 2(e). This effect can be explained by examining the laser field structure on the shoulders of the focus. The panels in the third row of Fig. 1 show the electric field time evolution at the focus center (solid lines) and at a lateral distance of $4\lambda$, where the angular momentum density is the highest (see plots in the second row of Fig. 1). Comparison of panels Fig. 1(h) and Fig. 1(i) shows that while the electric field at the focus center (solid lines) is almost the same for the cases of a single CP and of four LP pulses, the field at the focus periphery (dash-dotted lines) appears at its interference maxima higher for the four beam case. This interference effect, efficient enough taking into account that the magnetic field amplitude scales approximately as $a_{0}^{4}$ [20], makes the IFE stronger at larger distances from the beam axis. At the same time, although the electric field amplitude at the periphery keeps growing with the angle increasing (see dash-dotted lines in panels Fig. 1(j) and Fig. 1(k)), the angular momentum density drops down there. Therefore, though $a_{0}$ is sufficiently large out of the axis for bigger angles, the field polarization state is far from circular, which suppresses the angular momentum transfer at large distances. As a result, for larger angles the part of the focus where the IFE remains efficiently induced, shrinks down to the small vicinity of the x-axis.
  5. 5.
    In panels (g-i) of Fig. 2 we show results obtained for the same beam parameters as used for (e,b,d) but with the quantum recoil effect semi-classically accounted for through the factor $g(\chi)$ (6). Quantum corrections considerably damp the radiation reaction force leading to a visible decrease both in the magnetic field amplitude and in the size of space where the magnetic field is observed. The reduction factor is more than two (see also Fig. 2 in [22]). However, the spatial structure of the generated field and its dependence on the angle between the beams remain similar to that observed in the classical simulations. Note that the account of recoil through the semiclassical $g$-factor may lead to an overestimation of quantum effects compared to the case in which they are treated stochastically [45, 46] so that the magnetic field amplitudes in the counterpart panels (e)-(g), (b)-(h) and (d)-(i) should be considered as the upper and the lower bounds to the field, which could potentially be predicted in a full quantum calculation with stochastic radiation events and other quantum effects incorporated. The field distribution in Fig. 2(i) looks qualitatively different from that shown in all other panes. Excitation of a relatively strong field at $2\lambda\leq x\leq 5\lambda$ can be attributed to a surface mechanism not connected with the RRF [47]. This mechanism becomes more efficient with the angle between the pulse propagation and the vector normal to the surface increasing. Discussed contribution is present in panel (d) too, in the same space region, but there it remains much less visible because of a stronger contribution from the IFE.
@@
Secondly, for multi-PW laser sources with sub-kJ pulse energies and low repetition rates, it is hard to expect the beams to be well synchronized in phase and amplitude. For such laser systems under development, the option of fully incoherent beams is also considered as one of realistic scenarios. Thus it is of high relevance for the future experiments to make a step out of the idealistic assumption of fully coherent in-phase beams and to study sensibility of the effect to fluctuations of the relative phases and amplitudes of the beams. Fig. 3 presents the magnetic field distribution $B_{x}(x,r)/B_{0}$ recorded at $t=35T_{L}$ for the 4-beam case with $2\theta=5^{\rm o}$ and $2\theta=10^{\rm o}$. These cases were chosen basing on the analysis of the distributions present in Fig. 2. Panels (a-f) correspond to regular phase shifts between the pairs of beams with orthogonal polarizations, i.e. between the two blue and the two red beams in Fig.1(b), while panels (g) and (h) show the distributions generated by the beams with randomly shifted phases and amplitudes, correspondingly.
@@
Figure 3: Magnetic field distribution for the 4-pulse scheme with $2\theta=5^{\rm o}$ (a,c,e) and $2\theta=10^{\rm o}$ (b,d,f) calculated at $t=35T_{L}$ for the additional to $\pi/2$ phase shift $\Delta\varphi=0.0$ (a,b), $\Delta\varphi=0.2\pi$ (c,d) and $\Delta\varphi=0.4\pi$ (e,f). The field amplitude $a_{0}=250$ in each beam corresponds to the peak intensity $J\approx 1.4\cdot 10^{23}{\rm W}/{\rm cm}^{2}$ and power $P\approx 20$PW per beam. Panel (g) shows the distribution resulting from random relative phase shifts $0.075\pi,-0.018\pi,0.05\pi$; panel (h) – the distribution for randomly chosen amplitudes $a_{0}=184,242,162$ and $239$. The lines of constant magnetic field $B_{x}=[-15,-10,-5]B_{0}$ are displayed white.
@@
As expected, the efficiency of the magnetic field generation drops down with the phase shift increasing. This suppression can be qualitatively demonstrated by plotting contours of the constant magnetic field for different phase shifts. The result is shown in Fig. 4. It is important to note that the effect survives in a broad interval of relative phases. Comparison of panels (c), (d) (e) and (f) in Fig. 3 shows that the magnetic field remains visible at relative phases filling more than 50% of the interval $(0,\pi/2)$. The distributions in Figs.3 and 4 were calculated for higher field amplitudes $a_{0}=250$ and $a_{0}=500$. In the latter case, the 2-beam scheme was considered. A comparison of results presented in these Figures shows, in particular, that both the schemes are, on the qualitative level, equally sensitive to fluctuations of relative phases of the beams. In experiment, both the relative phases and amplitudes of the beams will be difficult to control. Panels (g) and (h) of Fig. 3 show the distributions for randomly chosen phase shifts and amplitudes. These results also suggest that the scheme is acceptably stable with respect to these parametric uncertainties.
@@
Thirdly, the degree of focusing also should play a certain role in the magnetic field excitation at the fixed pulse energy and power. We examined the sensitivity of the scheme to the tightness of the laser focus. The one used in Figs.2-4 with $r_{0}=3.8\lambda$ can be reduced up to $r_{0}=2\lambda$ keeping it still realistic for the state-of-the art focusing systems applied in PW lasers, see, e.g. [24]. Our simulations made with such tightly focused pulses showed a higher magnetic field or, equivalently, a lower power threshold per single beam. However, for the tighter focusing and non-vanishing angles of convergence $\theta$, another mechanism of the magnetic field excitation starts playing an important role. The contribution of this mechanism is clearly visible in Fig.2(i) as a magnetic field located at $2\lambda\leq x\leq 5\lambda$. It stems not from the effect of radiation friction, but from that of a sharp plasma edge, which leads to additional angular momentum transfer from the laser field to the plasma electrons at the laser incidence not normal to the plasma surface [47]. In the forthcoming publication we will examine this mechanism and its relative significance with respect to the angular momentum transfer induced by radiation friction.
@@ IV Conclusion
In conclusion, we have studied the possibility to reach the radiation-dominated regime of the laser plasma interaction and, in particular, to observe the Inverse Faraday Effect driven by radiation reaction in a multi-beam scheme. Our numerical results show that application of four beams partially synchronized in phase and amplitude lowers the single beam power required for reliable detection of the quasi-static longitudinal magnetic field excited through the IFE to 15 PW, which looks a realistic value for several forthcoming laser facilities. Most importantly, the effect is shown rather robust with respect to fluctuations of relative phases and amplitudes. This enhances the possibility of the observation at low repetition rates typical for multi-PW laser sources. The effect of quantum recoil of emitted photons reduces the radiation reaction force and results in a significant suppression of the excited quasi-static magnetic field, which however remains clearly visible for all considered sets of parameters of the 4-beam scheme.
An important feature observed in our simulations is the relatively low decrease of the effect with angle between the beams increasing. The latter leads to the following changes in the electromagnetic field in the focal area: (i) the total angular momentum carried by the field decreases (see Fig. 1(c)) – this is a global effect, and (ii) the spatial structure of the field deviates from that of a CP electromagnetic wave – a local effect. Our simulations show that the local field structure appears more important for the magnetic field excitation than the integrated angular momentum value. This can be clearly seen from comparison of e.g. Fig. 1(e) and Fig. 1(g) for $2\theta=5^{o}$ and $2\theta=20^{o},$ correspondingly. The value of $M_{x}$ changes the sign between these two points and even more, the absolute value of (negative) $M_{x}$ is higher for $2\theta=20^{o}$ than for $2\theta=5^{o}.$ However, some magnetic field of the same sign is still present in Fig. 2(f) and in Fig. 2(i). It is tightly localized near the axis $r=0,$ which agrees well with the fact that the angular momentum density remains positive in Fig. 1(g) close to the axis, too. Thus the IFE induced by radiation friction is rather sensitive to the angular momentum density than to its integrated value. This allows us to estimate the effective volume where a strong magnetic field can be created.
At the future XCELS facility [28], installations with up to 12 beams are considered. On one side, increasing the number of beams will obviously lower further the power threshold required for IFE observation. On the other side, taking into account the constraints on the angle $\theta$ and effects of the relative phase shifts we examined in this paper, application of $\sim 10$ or more beams may appear highly problematic from the technical viewpoint. We conclude that the four beam scheme at the 10–15PW level is optimal for the search of the giant magnetic field excitation in the radiation-dominated regime.